OpenStudy (anonymous):

solve for m s=4m-3n/21

OpenStudy (anonymous):

s = 4m-3n/21 please first clear my doubt is it s =(4m-3n)/21 or s = 4m - (3n/21)

OpenStudy (anonymous):

21s=4m-3n 21s+3n=4m m=21s+3n/(4)

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Can anyone tell me about the minimum value of m?

OpenStudy (anonymous):

i did right it is all right no mistake has been done !!

OpenStudy (anonymous):

i am confused because i am getting a different answer

OpenStudy (anonymous):

ok see here (4m-3n)/21 = s 4m-3n = 21s 4m = 21s + 3n m = (21s + 3n)/4

OpenStudy (anonymous):

this is what i got 7s+n/28

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

No it would be 3(7s + n)/4

OpenStudy (anonymous):

ok